# YOLO-LLTS: Real-Time Low-Light Traffic Sign Detection via Prior-Guided Enhancement and Multi-Branch Feature Interaction

### v2

@@ Abstract
Detecting traffic signs effectively under low-light conditions remains a significant challenge. To address this issue, we propose YOLO-LLTS, an end-to-end real-time traffic sign detection algorithm specifically designed for low-light environments. Firstly, we introduce the High-Resolution Feature Map for Small Object Detection (HRFM-TOD) module to address indistinct small-object features in low-light scenarios. By leveraging high-resolution feature maps, HRFM-TOD effectively mitigates the feature dilution problem encountered in conventional PANet frameworks, thereby enhancing both detection accuracy and inference speed. Secondly, we develop the Multi-branch Feature Interaction Attention (MFIA) module, which facilitates deep feature interaction across multiple receptive fields in both channel and spatial dimensions, significantly improving the model’s information extraction capabilities. Finally, we propose the Prior-Guided Enhancement Module (PGFE) to tackle common image quality challenges in low-light environments, such as noise, low contrast, and blurriness. This module employs prior knowledge to enrich image details and enhance visibility, substantially boosting detection performance. To support this research, we construct a novel dataset, the Chinese Nighttime Traffic Sign Sample Set (CNTSSS), covering diverse nighttime scenarios, including urban, highway, and rural environments under varying weather conditions. Experimental evaluations demonstrate that YOLO-LLTS achieves state-of-the-art performance, outperforming the previous best methods by 2.7% mAP50 and 1.6% mAP50:95 on TT100K-night, 1.3% mAP50 and 1.9% mAP50:95 on CNTSSS, and achieving superior results on the CCTSDB2021 dataset. Moreover, deployment experiments on edge devices confirm the real-time applicability and effectiveness of our proposed approach.
@@ I Introduction
Traffic sign detection plays a crucial role in advanced driver-assistance systems (ADAS) and autonomous vehicles, ensuring road safety and assisting in navigation. Existing object detection methods have demonstrated strong capabilities in accurately detecting various traffic elements, including pedestrians [1, 2, 3, 4], vehicles [5, 6, 7, 8, 9], and traffic lights [10, 11, 12]. However, the detection of traffic signs remains a considerable challenge, primarily due to their small size and the complexity involved in distinguishing them from other objects in the scene, making this task particularly difficult in real-world scenarios. In high-resolution images with dimensions of 2048$\times$2048 pixels, a sign may occupy only a small area of approximately 30$\times$30 pixels. Due to the extremely low resolution and limited information, significant efforts have been made in recent years to enhance the performance of small object detection. Existing traffic sign detection algorithms [13, 14, 15] have been improved to address the characteristics of small targets, enabling effective detection of traffic signs during the daytime.
@@
Figure 2: Application Scenarios of Traffic Sign Detection in advanced driver-assistance systems [16] [17] [18].
@@
To address the dual challenges of small targets with low resolution and low visibility in low-light environments, we have designed an end-to-end traffic sign recognition algorithm specifically for low-light conditions. As shown in Fig. 2, after the camera captures the traffic signs, the vehicle equipped with an assisted driving system performs calculations through a mobile edge computing device. The system uses our model YOLO-LLTS to detect the traffic signs, ensuring driving safety [23]. YOLO-LLTS improved YOLOv8 by utilizing high-resolution images to extract clearer features, and it also designed a new multi-branch features interaction attention module to fuse features from different receptive fields. Additionally, we have developed a prior-providing module that not only enhances the image but also supplements its details. Our algorithm effectively mitigates the poor performance of existing methods under low-light conditions, thereby improving the safety of autonomous driving systems. Furthermore, to address the lack of low-light scene data in existing traffic sign datasets, we have constructed a multi-scene Chinese traffic sign dataset ranging from dusk to deep night, providing a foundational resource for industry research in this area.
@@
    High-Resolution Feature Map for Small Object Detection (HRFM-TOD): To address the challenge of indistinct features for small objects in low-light conditions, we propose the HRFM-TOD module, which utilizes high-resolution feature maps for detection. This module effectively alleviates the feature dilution problem encountered in traditional PANet frameworks when detecting small targets, thus improving both detection accuracy and inference speed.
@@
    Prior-Guided Enhancement Module (PGFE): To overcome common image quality issues such as noise, reduced contrast, and blurriness inherent in low-light environments, we propose the PGFE module. This module utilizes prior knowledge to enhance images and supplement image details, significantly boosting detection performance under challenging low-light conditions.
@@ II-A Low-Light Image Enhancement Methods (LLIE Methods)
LLIE based on Machine learning methods mainly consist of two categories: supervised learning and unsupervised learning.
Supervised learning approaches generally rely on extensive datasets of low-light images as well as their paired images under normal-light conditions to facilitate effective training. For example, LLNet [30], as the first deep learning method for low light image enhancement (LLIE) that introduced an end-to-end network, was trained on simulated data with random gamma correction. Furthermore, Wei et al. [31] creatively combined Retinex theory with convolutional neural networks, dividing the network into modules of decomposition, adjustment and reconstruction, and trained it using their self-built LOL dataset. The performance of these methods relies heavily on the quality and diversity of the paired training datasets.
Unsupervised learning methods focus on enhancing low-light images without paired training data. For example, Zero-DCE [32] approached image enhancement as a task of estimating image-specific curves using a deep network, driving the network’s learning process through a series of meticulously crafted loss functions to achieve enhancement without the need for paired data. EnlightenGAN [33] used an attention mechanism-based U-Net [34] as a generator and employed the method of GAN to perform image enhancement tasks without paired training data. Cui et al. [35] proposed the Illumination Adaptive Transformer (IAT) model, which uses attention mechanisms to adjust parameters related to the image signal processor (ISP), effectively enhancing target under various lighting conditions. These methods demonstrate the potential for unsupervised learning in LLIE and demonstrate the flexibility and adaptability of deep learning models to diverse lighting conditions.
@@ III Methods
In this section, we provide a detailed introduction to our dataset and modules. Firstly, we describe the foundational aspects and detailed information of the dataset, which is intended to support further research in this area. Secondly, we introduce the design of the HRFM-TOD module, which alleviates the issue of blurred features of small objects in low-light conditions. Thirdly, we explain the design of the MFIA module, which enhances the model’s ability to perform feature interactions across multiple receptive fields. Finally, we present the PGFE module, which utilizes prior knowledge to improve the quality of low-light images.
@@ III-B High-Resolution Feature Map for Small Object Detection
Figure 6: Framework overview of our model YOLO-LLTS.
@@
To address this issue, We introduce the HRFM-TOD module to preserve detailed information in images while simultaneously reducing the network’s computational load. As shown in Fig. 6, deep multi-scale features are uniformly adjusted to the high-resolution image size $\times 160\times$128 through 1$\times$1 convolution and Bilinear upsampling. The formula is as follows:
@@
In the feature extraction process, we adjust the receptive field obtained from the backbone using the computational method proposed in [15]. As a result, the receptive field received by the HRFM-TOD module varies with the size of the target.
@@
Figure 7: Overview of Multi-branch Features Interaction Attention Module (MFIAi). MFIA is applied to $F_{i}$($i=1,2,3,4$) four times throughout the network. The illustration shows one MFIAi module along with its corresponding output feature $F_{i}^{\text{mfia}}$.
@@ III-D Prior-Guided Feature Enhancement Module
Figure 8: Overview of Prior-Guided Feature Enhancement Module (PGFE).
@@
To address this challenge, we propose the Prior-Guided Feature Enhancement (PGFE) Module, which replaces the original P0 layer of YOLOv8. This module transforms RGB images from 3 channels to 64 channels to enhance low-light images and provide prior knowledge for the following detection. The module consists of two main components: the Prior-Guided Enhancement (PGE) Module and the Detail Texture Recovery (DTR) Module.
@@ III-D1 Prior-Guided Enhancement Module
where $u_{t}$ represents the output parameters at the $t$-th stage ($t=0,…,T$), and $G(\cdot)$denotes the residual operation. $u_{1}$ represents the input feature, while $u_{T}$ represents the output feature. The output of $L(\cdot)$ is fed into $C(\cdot)$ for contrast enhancement, and then into $E(\cdot)$for edge enhancement. The principles of both operations are similar. $C(\cdot)$ enhances the contrast by amplifying the differences between different pixel values, and this can be represented by the following formula:
@@ IV-A3 CCTSDB2021
  Method Venue Input Size Precision(%) Recall(%) F1 mAP50(%) mAP50:95(%) Param(M) FPS GPU YOLOv5-L[62] 2020 640$\times$640 70.6 57.2 63.2 62.4 44.4 46.1 99.0 RTX 4090 YOLOv6-L[63] CVPR2022 640$\times$640 65.7 52.8 58.5 57.4 40.9 59.5 43.0 RTX 4090 GOLD-YOLO-L[64] NeurIPS2023 640$\times$640 71.3 54.7 61.9 60.5 43.1 75.0 55.2 RTX 4090 YOLOv8-L[65] 2023 640$\times$640 72.1 59.3 65.1 65.2 46.3 43.6 107.5 RTX 4090 YOLOv9-C[66] CVPR2024 640$\times$640 71.3 59.7 65.0 64.6 45.5 50.7 80.6 RTX 4090 YOLOv10-L[67] NeurIPS2024 640$\times$640 68.4 57.4 62.4 61.4 43.8 25.7 72.5 RTX 4090 Zhang et al. [21] TETCI2024 640$\times$640 17.3 18.9 18.1 21.9 - 20.8 105.1 RTX 4090 MIAF-net [14] TIM2024 640$\times$640 69.6 57.7 63.1 60.4 41.2 34.4 102.0 RTX 4090 YOLOv11-L[68] 2024 640$\times$640 69.5 57.3 62.8 62.5 44.1 25.3 73.0 RTX 4090 YOLO-TS[15] arXiv2024 640$\times$640 75.2 63.2 68.7 68.7 48.4 11.1 109.9 RTX 4090 YOLOv12-L[69] arXiv2025 640$\times$640 72.6 54.5 62.3 61.4 43.8 26.4 46.3 RTX 4090 YOLO-LLTS(ours) - 640$\times$640 77.2(+2.0$\uparrow$ 64.4(+1.2$\uparrow$ 70.2(+1.5$\uparrow$ 71.4(+2.7$\uparrow$ 50.0(+1.6$\uparrow$ 9.9 83.3 RTX 4090
@@
  Method Venue Input Size Precision(%) Recall(%) F1 mAP50(%) mAP50:95(%) Param(M) FPS GPU YOLOv5-L[62] 2020 640$\times$640 83.7 66.7 74.2 75.0 53.4 46.1 97.1 RTX 4090 YOLOv6-L[63] CVPR2022 640$\times$640 86.4 73.1 79.2 79.8 55.9 59.5 45.2 RTX 4090 GOLD-YOLO-L[64] NeurIPS2023 640$\times$640 86.0 73.9 79.5 79.9 56.2 75.0 42.9 RTX 4090 YOLOv8-L[65] 2023 640$\times$640 85.1 67.7 75.4 75.1 53.3 43.6 100.0 RTX 4090 YOLOv9-C[66] CVPR2024 640$\times$640 84.1 66.2 74.1 74.7 52.8 50.7 80.0 RTX 4090 YOLOv10-L[67] NeurIPS2024 640$\times$640 83.9 63.8 72.5 72.6 52.1 25.7 78.1 RTX 4090 Zhang et al. [21] TETCI2024 640$\times$640 42.7 44.6 43.6 48.9 - 20.8 110.8 RTX 4090 MIAF-net [14] TIM2024 640$\times$640 84.1 69.0 75.8 76.2 52.6 34.2 158.7 RTX 4090 YOLOv11-L[68] 2024 640$\times$640 80.4 66.9 73.0 73.2 51.9 25.3 65.4 RTX 4090 YOLO-TS[15] arXiv2024 640$\times$640 87.3 70.9 78.2 79.6 58.2 15.1 117.6 RTX 4090 YOLOv12-L[69] arXiv2025 640$\times$640 85.0 62.6 72.1 72.8 51.7 26.3 46.1 RTX 4090 YOLO-LLTS(ours) - 640$\times$640 88.3(+1.0$\uparrow$ 74.9(+1.0$\uparrow$ 81.0(+1.5$\uparrow$ 81.2(+1.3$\uparrow$ 60.1(+1.9$\uparrow$ 13.9 82.0 RTX 4090
@@
  Method Venue Input Size Precision(%) Recall(%) F1 mAP50(%) mAP50:95(%) Param(M) FPS GPU YOLOv5-L[62] 2020 640$\times$640 85.4 72.2 78.2 78.6 50.2 46.1 88.5 RTX 4090 YOLOv6-L[63] CVPR2022 640$\times$640 87.4 76.5 81.6 82.0 52.6 59.5 52.2 RTX 4090 GOLD-YOLO-L[64] NeurIPS2023 640$\times$640 83.8 77.0 80.3 80.6 51.2 75.0 63.3 RTX 4090 YOLOv8-L[65] 2023 640$\times$640 84.4 74.4 79.1 80.5 53.0 43.6 91.7 RTX 4090 YOLOv9-C [66] CVPR2024 640$\times$640 84.8 76.0 80.2 81.9 53.7 50.7 42.4 RTX 4090 YOLOv10-L[67] NeurIPS2024 640$\times$640 83.9 73.2 78.2 78.7 51.4 25.7 68.0 RTX 4090 Zhang et al. [21] TETCI2024 640$\times$640 82.7 80.7 81.6 78.1 - 20.8 101.1 RTX 4090 MIAF-net [14] TIM2024 640$\times$640 60.1 49.6 54.5 52.7 36.2 34.4 151.5 RTX 4090 YOLOv11-L[68] 2024 640$\times$640 85.8 74.2 79.6 81.2 54.1 25.3 70.4 RTX 4090 YOLO-TS[15] arXiv2024 640$\times$640 88.1 80.8 84.3 86.0 57.2 12.9 138.9 RTX 4090 YOLOv12-L[69] arXiv2025 640$\times$640 84.9 74.2 79.2 82.8 55.7 26.3 44.8 RTX 4090 YOLO-LLTS(ours) - 640$\times$640 88.8(+0.7$\uparrow$ 81.1(+0.3) 84.8(+0.5) 87.8(+1.8$\uparrow$ 57.5(+0.3$\uparrow$ 10.2 93.6 RTX 4090

### v3

@@ Abstract
Traffic sign detection is essential for autonomous driving and Advanced Driver Assistance Systems (ADAS). However, existing methods struggle with low-light conditions due to issues like indistinct small-object features, limited feature interaction, and poor image quality, which degrade detection accuracy and speed. To address this issue, we propose YOLO-LLTS, an end-to-end real-time traffic sign detection algorithm specifically designed for low-light environments. YOLO-LLTS introduces three main contributions: the High-Resolution Feature Map for Small Object Detection (HRFM-SOD) module to enhance small-object detection by mitigating feature dilution; the Multi-branch Feature Interaction Attention (MFIA) module to improve information extraction through multi-scale features interaction; and the Prior-Guided Feature Enhancement Module (PGFE) to enhance image quality by addressing noise, low contrast, and blurriness. Additionally, we construct a novel dataset, the Chinese Nighttime Traffic Sign Sample Set (CNTSSS), covering diverse nighttime scenarios. Experiments show that YOLO-LLTS achieves state-of-the-art performance, outperforming previous best methods by 2.7% mAP50 and 1.6% mAP50:95 on TT100K-night, 1.3% mAP50 and 1.9% mAP50:95 on CNTSSS, 7.5% mAP50 and 9.8% mAP50:95 on GTSDB-night, and superior results on CCTSDB2021. Deployment on edge devices confirms its real-time applicability and effectiveness.
@@ I Introduction
Traffic sign detection plays a crucial role in advanced driver-assistance systems (ADAS) and autonomous driving perception technology. ADAS systems and autonomous vehicles equipped with traffic sign recognition algorithms are able to obtain real-time traffic regulations and road information, thereby making accurate driving decisions to ensure road safety. Existing object detection methods have demonstrated strong capabilities in accurately detecting various traffic elements, including pedestrians [1, 2, 3, 4], vehicles [5, 6, 7, 8, 9], and traffic lights [10, 11, 12]. However, the detection of traffic signs remains a considerable challenge, primarily due to their small size and the complexity involved in distinguishing them from other objects in the scene, making this task particularly difficult in real-world scenarios. In high-resolution images with dimensions of 2048$\times$2048 pixels, a sign may occupy only a small area of approximately 30$\times$30 pixels. Due to the extremely low resolution and limited information, significant efforts have been made in recent years to enhance the performance of small object detection. Existing traffic sign detection algorithms [13, 14, 15] have been improved to address the characteristics of small targets, enabling effective detection of traffic signs during the daytime.
@@
Figure 2: Application Scenarios of Traffic Sign Detection in advanced driver-assistance systems [16, 17, 18].
@@
To address the dual challenges of small targets with low resolution and low visibility in low-light environments, we have designed an end-to-end traffic sign recognition algorithm specifically for low-light conditions. As shown in Fig. 2, low-light conditions can impair the driver’s visibility, leading to the neglect of traffic signs and potential violations of traffic regulations. The ADAS, equipped with our YOLO-LLTS model, alerts the driver to take corrective actions or directly controls the vehicle upon detecting traffic signs. For example, when the system detects a speed limit sign, it automatically adjusts the vehicle’s speed to comply with the road’s speed limit. Similarly, if a stop sign is detected, the system either alerts the driver or automatically applies the brakes. YOLO-LLTS improved YOLOv8 by utilizing high-resolution images to extract clearer features, and it also designed a new multi-branch features interaction attention module to fuse features from different receptive fields. Additionally, we have developed a prior-providing module that not only enhances the image but also supplements its details. Our algorithm effectively mitigates the poor performance of existing methods under low-light conditions, thereby improving the safety of autonomous driving systems. Furthermore, to address the lack of low-light scene data in existing traffic sign datasets, we have constructed a multi-scene Chinese traffic sign dataset ranging from dusk to deep night, providing a foundational resource for industry research in this area.
@@
    High-Resolution Feature Map for Small Object Detection (HRFM-SOD): To address the challenge of indistinct features for small objects in low-light conditions, we propose the HRFM-SOD module, which utilizes high-resolution feature maps for detection. This module effectively alleviates the feature dilution problem encountered in traditional PANet frameworks when detecting small targets, thus improving both detection accuracy and inference speed.
@@
    Prior-Guided Feature Enhancement Module (PGFE): To overcome common image quality issues such as noise, reduced contrast, and blurriness inherent in low-light environments, we propose the PGFE module. This module utilizes prior knowledge to enhance images and supplement image details, significantly boosting detection performance under challenging low-light conditions.
The remaining parts of this article are organized as follows. Section II reviews the related works on low-light image enhancement methods, target detection methods in complex scenarios, and small target detection methods. Section III describes the details of the CNTSSS dataset we constructed and the low-light traffic sign detection algorithm YOLO-LLTS. Section IV presents the experimental analysis and results. Finally, Section V concludes the paper and discusses future work.
@@ II-A Low-Light Image Enhancement Methods (LLIE Methods)
LLIE based on Machine learning methods mainly consist of two categories: supervised learning and unsupervised learning. Supervised learning approaches generally rely on extensive datasets of low-light images as well as their paired images under normal-light conditions to facilitate effective training. For example, LLNet [29], as the first deep learning method for low light image enhancement (LLIE) that introduced an end-to-end network, was trained on simulated data with random gamma correction. Furthermore, Wei et al. [30] creatively combined Retinex theory with convolutional neural networks, dividing the network into modules of decomposition, adjustment and reconstruction, and trained it using their self-built LOL dataset. The performance of these methods relies heavily on the quality and diversity of the paired training datasets. Unsupervised learning methods focus on enhancing low-light images without paired training data. For example, Zero-DCE [31] approached image enhancement as a task of estimating image-specific curves using a deep network, driving the network’s learning process through a series of meticulously crafted loss functions to achieve enhancement without the need for paired data. EnlightenGAN [32] used an attention mechanism-based U-Net [33] as a generator and employed the method of GAN to perform image enhancement tasks without paired training data. Cui et al. [34] proposed the Illumination Adaptive Transformer (IAT) model, which uses attention mechanisms to adjust parameters related to the image signal processor (ISP), effectively enhancing targets under various lighting conditions. These methods demonstrate the potential for unsupervised learning in LLIE and demonstrate the flexibility and adaptability of deep learning models to diverse lighting conditions.
@@ III Methods
In this section, we provide a detailed introduction to our dataset and modules. Firstly, we describe the foundational aspects and detailed information of the dataset, which is intended to support further research in this area. Secondly, we introduce the design of the HRFM-SOD module, which alleviates the issue of blurred features of small objects in low-light conditions. Thirdly, we explain the design of the MFIA module, which enhances the model’s ability to perform feature interactions across multiple receptive fields. Finally, we present the PGFE module, which utilizes prior knowledge to improve the quality of low-light images.
@@ III-B High-Resolution Feature Map for Small Object Detection
Figure 6: Framework overview of our model YOLO-LLTS. The structure of our YOLO-LLTS model is divided into three components: Backbone, Neck, and Head. The Backbone is responsible for feature extraction, the Neck handles multi-scale feature fusion, and the Head is responsible for final object detection and classification. The main contributions focus on the PGFE module in the Backbone, the introduction of HRFM-SOD as a new Neck, and the MFIA module within the HRFM-SOD.
Figure 7: Some module design structures are derived from YOLOv8. (a), (b), and (c) represent the CBS module, SPPF module, and C2f module, respectively.
@@
To address this issue, We introduce the HRFM-SOD module to preserve detailed information in images while simultaneously reducing the network’s computational load. As shown in Fig. 6, deep multi-scale features are uniformly adjusted to the high-resolution image size $\times 160\times$128 through 1$\times$1 convolution and Bilinear upsampling. The formula is as follows:
@@
The design details of the module architecture of YOLOv8 are shown in Fig. 7. In the feature extraction process, we adjust the receptive field obtained from the backbone using the computational method proposed in [15]. As a result, the receptive field received by the HRFM-SOD module varies with the size of the target.
@@
Figure 8: Framework overview of Multi-branch Features Interaction Attention Module (MFIA). (a) shows the overall architecture of the MFIAi module. MFIA is applied to $F_{i}$($i=1,2,3,4$) four times throughout the network. The illustration shows one MFIAi module along with its corresponding output feature $F_{i}^{\text{mfia}}$. The MFIA is divided into two parts: the MFCA and MFSA. (b) shows the CAM, which is used twice within the MFCA. (c) illustrates the SAM in the MFSA.
@@ III-C2 Multi-branch Features Spatial Attention Module
Figure 9: Overview of Prior-Guided Feature Enhancement Module (PGFE). The PGFE module is divided into two parts: the Prior-Guided Enhancement Module (PGE) and the Detail Texture Recovery Module (DTR).
@@ III-D Prior-Guided Feature Enhancement Module
To address this challenge, we propose the Prior-Guided Feature Enhancement (PGFE) Module. This module transforms RGB images from 3 channels to 64 channels to enhance low-light images and provide prior knowledge for the following detection. As shown in Fig. 9, the module consists of two main components: the Prior-Guided Enhancement (PGE) Module and the Detail Texture Recovery (DTR) Module.
Figure 10: (a) shows the structure of L(u) in the PGE, which enhances features while improving image brightness. The pink dashed box represents G(·) from Eq. 7, denoting the residual operation. (b) shows the structure of the CBR, which is used multiple times within the residual operation in L(u).
@@ III-D1 Prior-Guided Enhancement Module
where $u_{t}$ represents the output parameters at the $t$-th stage ($t=0,…,T$), and $G(\cdot)$ denotes the residual operation. $u_{1}$ represents the input feature, while $u_{T}$ represents the output feature.
The output of $L(\cdot)$ is fed into $C(\cdot)$ for contrast enhancement, and then into $E(\cdot)$for edge enhancement. The principles of both operations are similar. $C(\cdot)$ enhances the contrast by amplifying the differences between different pixel values, and this can be represented by the following formula:
@@
Figure 11: (a) shows the structure of the DTR in the PGFE, which is employed to supplement the detailed information lost due to feature enhancement. (b) shows the structure of the CDC in the DTR, where the CDC utilizes Depthwise Convolution to extract local features from the image.
@@ IV-A3 CCTSDB2021
IV-A4 GTSDB-night
The GTSDB dataset consists of 900 images, with 600 images allocated for training and 300 images for testing. This dataset covers 43 common traffic sign categories found in Germany. Each image has a resolution of $1360\times 800$ pixels and may contain one or more traffic signs, or none at all. Similarly, we use CycleGAN to generate nighttime images from daytime images in order to evaluate the model’s performance under low-light conditions.
@@ IV-A4 GTSDB-night
[…]
@@
[…]
@@
[…]
@@
TABLE IV: Performance comparison on GTSDB-night dataset; The first and second best results are indicated in blue and green
[…]